OpenStudy (anonymous):

x 1 2 3 4 f(x) 0 1 2 3 g(x) -1 -2 -4 -5 f'(x) 4 3 5 6 g'(x) -3 0 -2 1 compute (g of f)' (4)

OpenStudy (rulnick):

I hope I'm understanding what is meant by (g of f)'(4). Here goes: let h(x) = g(f(x)) h'(x) = g'(y) f'(x) evaluated at y=f(x) (chain rule) For x=4 this is h'(4) = g'(f(4)) f'(4) = g'(3) f'(4) = -2 (6) = -12

OpenStudy (anonymous):

thats right. thanks for the help

OpenStudy (rulnick):

oh, good. you're welcome.